OpenStudy (anonymous):

Gretchen worked in the ticket booth for 20 minutes. Kevin worked in the ticket booth 5 times longer than Gretchen. Which equation could be used to calculate the number of minutes that Kevin worked? Let the variable b stand for the number of minutes Kevin worked. A. 20 = b + 5 B 20 = b * 5 C b = 20 / 5 D b = 5 * 20

OpenStudy (anonymous):

Is B right thanks

OpenStudy (anonymous):

@AloneS

OpenStudy (anonymous):

@jayjaysimmer

OpenStudy (anonymous):

@pooja195

OpenStudy (anonymous):

Somebody...

OpenStudy (kkutie7):

G=20 K=5*G Or K=5*20

OpenStudy (kkutie7):

Are you having trouble?

OpenStudy (anonymous):

yes

OpenStudy (kkutie7):

Gretchen= 20 is a representation of how many min she worked. I use G to represent her. I used K to represent how long Kevin worked compared to Gretchen

OpenStudy (kkutie7):

Kevin worked five time the amount of Gretchen. So I wrote K=5*G From before we know G=20 Substitute 20 in for G and you get K=5*20

OpenStudy (kkutie7):

In your problem they uses b instead of K

OpenStudy (anonymous):

b= 5 x 20 D right thanks and sorry

OpenStudy (kkutie7):

Don't be sorry